Report of Activities on the RV Ron Brown CMarZ Cruise 06-03
15/16 April 2006

In the early morning of the 15th of April, there was bright sun coming through the portholes and mix of sun and clouds. The winds were from west southwest at about 25 kts. We were in the southern fringe of the gale that had been forecast for just north of our station #1 for today and there were lots of white caps on the starboard bow as we steamed south. Fortunately we were steaming away from the gale as we headed for Station #2 at 30N; 70W.

During the day, the main laboratory was the center of activity with the taxonomic specialists sorting through the samples collected at Station #1, identifying zooplankton species, and working with Team DNA to prepare the specimens for sequencing.

In mid-afternoon, we crossed paths with the lead boat in the Volvo Ocean Race. The leg five leader, "ABN AMRO ONE" was off to starboard about 4.5 nm and was headed for Baltimore. There are seven boats in the race. In addition to "ABN AMRO ONE", there is "movistar", "Pirates of the Caribbean", "Brasil 1", "Ericsson Racing Team", "ABN AMRO TWO", and "Brunel".

The station work began about 1800 with the deployment of the 1/4-m MOC. It did not go well. The same problem that occurred on the previous tow at the last station occurred on this one. About 250 m down and with 600 meters of wire out, the deck unit lost the underwater signal and it could not be re-established. There was no indication of a problem up to the time the system failed. After several hours of testing, two of the underwater units were unable to work on the CTD cable. In addition, one underwater unit suffered electronic damage as a result of the failure. Since these units worked well on the trawl wire, the suspicion was that the CTD cable was causing the problem. This ended our attempt to make a 1/4-m tow at this station.

The night dive that was scheduled for 2130 was also scrubbed because the winds (17 to 21 kts) and sea state were marginal at best. We then were going to do a 1-m MOCNESS early since the time was available and the deployment was scheduled for 2230. Leo Bercial, Chaolun Li, Paola Lona, and Joe Catron got the cod-end buckets on and set the net bar traps. Bruce, the Bosun, came out onto the deck at 2230 and started the operation by picking up the frame with the port crane to move it to the deck area underneath the A-frame. The MOC-1 is set up in a different place, forward of the MOCNESS-10. The tow cable got caught on the support stanchion just as the frame was being lifted rapidly and the cable was stressed to the breaking point. It took about an hour and a half to re-splice and water-proof the cable. About midnight on the 15th, the 1-m2 MOCNESS was ready to go back into the water and indeed it did.

During the wee hours of the 16th of April, the 1-m2 MOCNESS tow was successfully completed. While the samples were being processed and the net system being reset for another tow, 30-liter bottle casts to collect water for tintinnid analyses took place. This was followed by a pair of vertical Reeve net tows.

The second 1-m2 MOCNESS at station #2 started about 0730 on April 16th and was successfully completed by 1000. Again there was a flurry of activity as the samples came off the deck and went into the cold room. Those processing the samples were working quickly because the nets came on board just as the group was finishing up the preservation of the previous tow's samples.

Although a blue-water dive was planned for the late morning, wind (around 20 kts out of the west southwest) and sea conditions were still marginal, in spite of the fact that it was warm and sunny, and it was delayed until early afternoon. Then with conditions a bit worse, it was cancelled.

In the early afternoon, Erich Horgan and others on the afternoon watch worked on setting up the 10-m2 MOCNESS. A bracket with a new tab, made by the ship's engineers, was mounted on net bar 3 and then a cable with the swaged fitting was fixed through the tab. The nets were loaded and nets laid out ready to be deployed. During this tow, a series of test stops and haul backs were planned to make sure the winch could retrieved the wire and to allow the ship's engineers to make adjustments to improve the winch performance.

In spite of the preparation, the launch and towing of the MOC-10 proved to be an ordeal. The launch was for 1445. The first snag was with the figee fitting (cable termination). When the cable was hauled up tight in the over-boarding sheave, it was clear that the electrical cable was twisted in a way that had to be fixed. So we had to take the termination apart and reconfigure the electrical wire so that it was out of harms way. Then we put in the cod-end buckets and nets, which had been laid out so that the cod-ends could be easily deployed. In the process, a long rent was discovered in the net with 3mm mesh that is deployed open (net zero) for the trip to depth. So the launch was stopped while the net was sewn up. Starting the launch again, the tie-down straps were released, the tugger lines were slacked and unhooked, and then the winch wire was hauled in to raise the frame off the deck. Immediately it was noticed that the bar for net 1 was released from the toggle release. So the launch was stopped again, the air-tuggers hooked back onto the frame and the frame brought back down onto the stanchion. Then some agonizing time was spent trying to get the net bars up to the top of the frame so that the swaged fitting from net 1 could be inserted into the toggle and latched. This was made exceedingly difficult because the nets were streaming behind the ship and their drag was very hard to counter. Finally the fitting was secured. The tuggers were again released and the frame was lifted up and rolled down into the water, and the tow began.

During the down-leg of this tow, the wire went out at a steady 30 m per minute until reaching 4500 meters of wire out (MWO). Then the Chief Engineer had the winch stop and haul in about 250 m to test the winch's haul-in capability and to make some adjustments. After that it went down again. But the ship was having trouble with the wire tending to starboard of the stern because it was no longer towing into the wind. In order to get the ship pointed into the wind, the bridge decided to make a slow turn to port until the ship was again steaming into the wind. That maneuver was started when the wire was at 5000 MWO. The change in the course had a negative consequence. Towards the end of the turn, which started at 1906 and ended at 2005, the net began rapidly descending vertically, reaching speeds of 60 to 88 m/minute. Essentially the net was in free-fall. An increase in ship's speed was requested, but the time lag for an effect was long (on the order of minutes to ten minutes or more). Later when the net system was brought on board, a kink was in the wire about 5 meters above the cable termination. The kink probably occurred when the free-fall period stopped and tension was restored on the wire. Once the ship was headed into the wind, the wire streamed straight out off the stern and the net frame stabilized. There was one more test of the wire at 5500 MWO, where the winch was stopped and the wire hauled in a few hundred meters at around 15 m/min. This worked OK. Then the wire was paid out to 6000 MWO (the limit for this tow) and the ship's speed was reduced to allow the net down to the desired maximum sampling depth. However, this was done with limited success. The net reached a depth of 4315 m before it started to ride back up. And that is where net zero was closed and net 1 opened. Somewhat later the winch operator saw that the level-wind had malfunctioned and that there was a bad wrap on the winch. He had to stop and pay out wire to fix it. It turned out to be around 600 m. The rest of the haul went fairly smoothly, although there were additional problems with the level-wind.

The trawl came back on board about 0400 on the 17th of April. Several of the samples from great depths appeared to be contaminated by animals living closer to the sea surface. This is a problem that often goes unnoticed when towing opening/closing nets systems shallower than 1000 m where life forms are abundant. But when sampling the bathypelagic realm where most species occur in very low abundance, even small amounts of contamination from the nets passing through the shallow waters and catching surface forms can be significant. Still the nets caught some important species of zooplankton and fish that had not been caught in previous tows. In addition, a bathypelagic fish was caught that is likely a new species (see below).

After the trawl was brought on board, two Reeve net and two surface ring net tows were conducted. Although a second trawl was scheduled for Station #2, it was cancelled because of the number of repairs needed to be made to the trawl and trawl wire, and also because the winds and seas were building. Winds in excess of 30 kts were forecast for the area later in the day. Around 0700 on 17 April, the ship started steaming for station #3 approximately 600 nm to the southeast [25N; 60W].

While the over-the-side work at Station #2 was trying, the sample collection has provided the biologists with more material from which to work and a number of new species were added to the list of those already identified. The following three brief reports provide an indication of the exciting work taking place on this cruise.

Tracey Sutton Fish Report: The second big MOCNESS tow revealed several exciting specimens including a possible new species of great swallower of the genus Pseudoscopelus. Our specimen, extremely large and in near perfect condition, is quickly differentiated from its sister species by a unique combination of jaw teeth, lateral line pores, and photophore series. Detailed examination will follow. Other catches of note included a second rare male anglerfish, a giant loosejaw dragonfish, and a juvenile oarfish, which as an adult can reach 11 m (in fact listed in Guinness Book of Records as world's longest bony fish).

Dhugal Lindsay Report: In the MOCNESS plankton net with the mouth area of one square meter we caught an amazing little squid that looks somewhat like a writhing strawberry (Figure 1). This little squid belongs to the genus Histioteuthis, the cock-eyed squids, and had obviously forayed up into surface waters for its nighttime feeding frenzy as it was caught between 50 and 100 m depth at around 03:30 in the morning on 16 April. Squid of this genus are covered in little photophores, or light-producing organs, which give the squid's reddish skin the pockmarked appearance of a strawberry. These squid have one eye larger than the other and it is supposed that one eye looks down into the murky depths while the other eye looks upwards for shadows against the downwelling light. Unfortunately this little beauty expired while we were sorting the fragile gelatinous forms such as ctenophores and other jellyfishes from the plankton samples so we were unable to do behavioral experiments on it. Perhaps one day, a crewed submersible or a remotely-operated vehicle will be able to observe this beautiful animal in its own environment and let us know just how this squid behaves in its native realm.

Figure 1. The cock-eyed squid caught in the 1-m2 MOCNESS.
(photo by Dhugal Lindsay)

Team DNA second report: Since the end of station 1 and our last report, Team DNA has been hard at work on two fronts. The humidity and high salt content in the air has caused problems with our sequencing reaction clean-up procedure leading to lots of fluorescent noise on the sequencer. The team designed a series of experiments to diagnose the problem and develop a modified clean-up protocol. Without betraying too many details, I am at liberty to say that the winning method involved the addition of something borrowed from the tables in the mess hall. On the other front, we have taken the time created by the sequencing slow-down to improve our cataloguing procedure, and bring the rest of our samples as far along as possible. Station 2 gave us a number of beautiful new specimens and intriguing questions to pursue with our molecular techniques. The sample count now stands at some 360 individuals comprised of around 170 species. Now that our throughput is back up to where we want it, processing will go much faster and we will have plenty of good data to crunch.